OpenStudy (legomyego180):

A 1000-gallon tank, initially full of water, develops a leak at the bottom. Given that 500 gallons of water leak out in the first 30 minutes, find the amount of water left in the tank t minutes after the leak develops given that the water drains off a rate proportional to the amount of water present.

OpenStudy (sdfgsdfgs):

calculus right? let Q be the quantity of water in tank dQ/dt will be the rate of change of water in tank (drain rate) given that the water drains off a rate proportional to the amount of water present dQ/dt = kQ, where k is a constant Integrate n use the given "500 gallons of water leak out in the first 30 minutes" to solve for k. Good luck!

OpenStudy (sshayer):

\[\frac{ dQ }{ dt }=kQ\] separating the variables and integrating \[\int\limits \frac{ dQ }{ Q }=\int\limits k~dt+c\] \[\ln Q=kt+c\] \[Q=e ^{kt+c}=e^c~e ^{kt}=a e ^{kt}\] when t=0,Q=1000 g \[1000=e^ce^0=e^c\] \[Q=1000~e ^{kt}\] when t=30 min. Q=1000-500=500g \[500=1000e^{30k}\] \[\frac{ 500 }{ 1000 }=e^{30k},\frac{ 1 }{ 2 }=e^{30k},30k=\ln \frac{ 1 }{ 2 }=\ln 1-\ln 2=0-\ln 2=-\ln 2\] \[k=\frac{ -\ln~2 }{ 30 }\] replace the value of k and findQ

OpenStudy (legomyego180):

Thanks for the help guys, sorry my computer died earlier and I didnt have the charger. Ended up getting the correct answer with your help so thank you